Give examples of two metrics on R ^2 which are not equivalent .substantiate your claim
1
Expert's answer
2021-07-16T13:28:31-0400

Solution:

The following metrics on "\mathbb{R}^{2}" are not equivalent to one another: the Euclidean metric d, the hub metric "\varrho_{h}" , and the discrete metric "\varrho_{\text {disc }}". These are defined as follows:


The euclidean metric on "\mathbb{R}^{2}" is defined by

"d(\underline{x}, \underline{y})=\sqrt{\left(x_{1}-y_{1}\right)^{2}+\left(x_{2}-y_{2}\right)^{2}},"

where "\underline{x}=\left(x_{1}, x_{2}\right)" and "\underline{y}=\left(y_{1}, y_{2}\right)"


Let "X=\mathbb{R}^{2}" . For "x=\left(x_{1}, x_{2}\right), y=\left(y_{1},y_{2}\right)" define "\varrho_{h}(x, y)" as follows. If x=y then "\varrho_{h}(x, y)=0" . If "x\ne y" then

"\varrho_{h}(x, y)=\sqrt{x_{1}^{2}+x_{2}^{2}}+\sqrt{y_{1}^{2}+y_{2}^{2}}"

The metric "\varrho_{h}" is called the hub metric on "\mathbb{R}^{2}" .


The discrete metric on an arbitrary set X is defined by

"d(x, y)=\left\{\begin{array}{lll}

0 & \text { if } & x=y \\

1 & \text { if } & x \neq y

\end{array}\right."


Every metric defines open balls, but even if metrics are equivalent their open balls may look very differently (compare e.g. open balls in "\mathbb{R}^{2}" taken with respect to d and "\left.\varrho_{\text {ort }}\right)" . It turns out, however, that each metric defines also a collection of so-called open sets, and that open sets defined by two metrics are the same precisely when these metrics are equivalent.